### v3

# A Short Character Sum in $\mathbb{F}_{p^{3}}$

Abstract.
We establish a new bound for short character sums in finite fields, particularly over two-dimensional grids in $\mathbb{F}_{p^{3}}$ and higher-dimensional lattices in $\mathbb{F}_{p^{d}}$, extending an earlier work of Mei-Chu Chang [8] on Burgess inequality in $\mathbb{F}_{p^{2}}$. In particular, we show that for intervals of size $p^{3/8+\varepsilon}$, the sum $\sum_{x,y}\chi(x+\omega y)$, with $\omega\in\mathbb{F}_{p^{3}}\setminus\mathbb{F}_{p}$, exhibits nontrivial cancellation uniformly in $\omega$. This is further generalized to codimension-one sublattices in $\mathbb{F}_{p^{d}}$, and applied to obtain an alternative estimate for character sums on binary cubic forms.
Key words and phrases:
Short Character Sum, Multiplicative Energy, Weil’s Bound, Binary Cubic Forms
2010 Mathematics Subject Classification:
11L26,11L40.
1. Introduction
Short-character sums are of importance in analytic number theory. By "Short character sum", we mean that the summation is taken over an interval $I$ of size smaller than $q^{\frac{1}{2}}$ where $q$ is the modulus of the character. In our article, characters are always multiplicative, and intervals are intervals of integers, more precisely $a,b]:=\{n\in\mathbb{Z}:a\leq n\leq b\}$. Standard applications of these sums involve determining the smallest quadratic non-residue $(\bmod{q})$ and the distribution of primitive elements in finite fields. We shall revisit some history in this connection. Burgess’ bound [3] on Character Sums $\sum_{x\in I}\chi(x)$ of Dirichlet character in prime fields $\mathbb{F}_{p}$ provide a nontrivial estimate for an interval $I\subset[1,p-1]$ of size $|I|>p^{\frac{1}{4}+\varepsilon}$ for any $\varepsilon>0.$ Burgess result, which improves the Polya-Vinogradov inequality if $|I|<p^{1/2}$, was a major breakthrough. Assuming the Generalized Riemann Hypothesis (G.R.H), however, one can already obtain a non-trivial upper bound as soon as $|I|>p^{\varepsilon}$, leaving significant space for improvement in the unconditional case. Later, Granville and Soundararajan [13] showed that to achieve a non-trivial bound, one must have $|I|\gg(\log q)^{A}$ for any fixed positive $A$. In a subsequent paper [14], they further established that a non-trivial upper bound exists even under a weaker hypothesis—namely, that “100%” of the zeros of $L(s,\chi)$ up to height $1/4$ lie on the critical line.
Davenport and Lewis [9] considered the character sum in $\mathbb{F}_{p^{d}}$ of the form of a linear form $F(x_{1},\ldots,x_{d})=x_{1}\omega_{1}+\ldots+x_{d}\omega_{d}$ for $\omega_{1},\ldots,w_{d}$ a fixed basis of $\mathbb{F}_{p^{d}}$. They proved
$\sum_{x_{1}\in I_{1},\ldots,x_{n}\in I_{d}}\chi(F(x_{1},\ldots,x_{n}))=O(|I_{1}|.\dots|I_{d}|.q^{-\delta(\varepsilon)})$
provided that $|I_{i}|=p^{\rho_{d}+\varepsilon}$ for all $i=1,\ldots,d$ where $\rho_{d}=\frac{1}{2}-\frac{1}{2(d+1)}$. In addition, if one assumes that $\omega_{1},\ldots,w_{d}$ is a certain special type of basis, stronger results have been obtained by Burgess [3] and Karacuba [17]. Without such special assumptions, Mei-Chu Chang [7] improved on the result of Davenport and Lewis (for $n\geq 5$) by obtaining a non-trivial upper bound for the size intervals $p^{\frac{2}{5}+\varepsilon}$. Furthermore, in dimension $d=2$, Chang [8] proved a non-trivial upper bound if $|I_{i}|=q^{1/4+\varepsilon}$ for any $\varepsilon>0$. Finally a generalization of Chang’s result to $\mathbb{F}_{p^{d}}$ and an improvement of Davenport and Lewis were established by Konyagin [18] for $|I_{i}|=q^{1/4+\varepsilon}$. In connection to these, we have our main theorem in $\mathbb{F}_{p^{3}}$.
Theorem 1.
Let $\varepsilon>0$. Then there exists a constant $\delta=\delta(\varepsilon)>0$ such that for all sufficiently large primes $p$, the following holds. If $\omega\in\mathbb{F}_{p^{3}}\setminus\mathbb{F}_{p}$ and $I,J\subset\mathbb{F}_{p}$ are intervals of length at least $p^{3/8+\varepsilon}$, then for every nontrivial multiplicative character $\chi$ of $\mathbb{F}_{p^{3}}$,
$\left|\sum_{x\in I}\sum_{y\in J}\chi(x+\omega y)\right|\;\;<\;\;p^{-\delta(\varepsilon)}\,|I|\,|J|.$
The importance of the above statement lies in it’s uniformness in $\omega$.
Remark 1: If we put $d=3$ and $\omega_{i}=\omega^{i-1}$ in the result of Davenport and Lewis then the length of the intervals becomes $\rho_{d}=\rho_{3}=3/8$. Unlike their theorem, our result allows one interval to be of smaller size.
Remark 2: For $\chi|{\mathbb{F}_{p}}\neq 1$ the theorem 1 follows from [12, Theorem]. However, our theorem also covers the case $\chi|_{\mathbb{F}_{p}}=1$ which was not encompassed in the previous results.
More generally, M.Chang [8, Theorem 10] established that if $\omega\in\mathbb{F}_{p^{d}}$ is a generator of $\mathbb{F}_{p}^{*}$ and if $I_{0},I_{1},\ldots,I_{d-1}$ are intervals of size $p^{\rho^{\prime}_{d}+\varepsilon}$, where $\rho^{\prime}_{d}=\frac{\sqrt{d^{2}+2d-7}+3-d}{8}$ then character sum over $\mathbb{F}_{p^{d}}$ is bounded by
$\sum_{x_{0}\in I_{0},x_{1}\in I_{1},\ldots,x_{d}\in I_{d}}\chi(x_{0}+x_{1}\omega+\dots+x_{d-1}\omega^{d-1})<p^{-\delta(\varepsilon)}|I_{0}||I_{1}|\ldots|I_{d-1}|$
where the above $\delta(\varepsilon)>0$ is independent of the choice of $\omega$. Now, it is easy to check $\rho^{\prime}_{d}<\rho_{d}$. So, it is an improvement over the size of the intervals mentioned by Devanport and Lewis [9]. Our second main theorem is an extension of the previous Theorem to a subset of the basis. Unlike M. C. Chang’s theorem over the complete basis of $\mathbb{F}_{p^{d}}$, we consider character sums evaluated over sublattices of co-dimension 1 in $\mathbb{F}_{p^{d}}$. We prove the following
Theorem 2.
Let $\chi$ be a nontrivial multiplicative character of $\mathbb{F}_{p^{d}}$and $\omega\in\mathbb{F}_{p^{d}}$ be a generator over $\mathbb{F}^{*}_{p}$. Given $\varepsilon>0$, there exists a $\delta(\varepsilon)$ such that if $J_{0},J_{1},\ldots,J_{d-2}$ be intervals of size $p^{\rho^{\prime\prime}_{d}+\varepsilon}$, where
$\rho^{\prime\prime}_{d}=\rho=\frac{1460-1000d+\sqrt{1000000d^{2}-1960000d+490000}}{960}$
Then
$\sum_{x_{0}\in I_{0},\ldots,x_{d-2}\in I_{d-2}}\chi(x_{0}+x_{1}\omega+x_{2}\omega^{2}+\ldots+x_{d-2}\omega^{d-2})<p^{-\delta(\varepsilon)}|J_{0}||J_{1}|\ldots|J_{d-2}|$
where, $\delta=\delta(\varepsilon)>0$ is independent of $\omega$.
The novelty of our results is that we consider character sums for sublattices of codimension 1. This has not been studied in earlier work.
Next, we shall apply this result to obtain a bound for the short character sum evaluated on homogeneous binary cubic forms over $\mathbb{F}_{p}$. It can be stated as follows:
Theorem 3.
Given $\varepsilon>0$, there is a $\delta(\varepsilon)>0$ such that the following holds: Let $\chi$ be a non-principle character in $\mathbb{F}_{p}$. Let $p$ be a large prime and $f(x,y)=x^{3}+ax^{2}y+bxy^{2}+cy^{3}$ which is non-degenerate over $\mathbb{F}_{p}$. Let $I,J\subseteq[1,p-1]$ be interval of size
$|I|,|J|=p^{\frac{3}{8}+\varepsilon}$
Then
$\left|\sum_{x\in I,y\in J}\chi(f(x,y))\right|<p^{-\delta(\varepsilon)}|I||J|.$
This estimate is uniform in $f$.
Let us recall some previous studies related to such sums. For $f(x,y)$ quadratic homogeneous polynomials, Burgess [6] proved the same type of estimate, where the interval length was $p^{\frac{1}{3}+\varepsilon}$ for large $p$. Chang [8] showed that if $f(x,y)$ is a quadratic form which is not a perfect square $\bmod{p}$, then there is a nontrivial upper bound if both intervals are of size $p^{\frac{1}{4}+\varepsilon}$. The above is a generalization of M.C. Chang’s result [8, Theorem 11] to cubic binary forms. Recently, Heath-Brown[15] established that Chang’s result continues to hold for primitive characters even if $q$ is any odd square-free modulli with $(q,det(f))=1$. In a previous joint work with Baier [1, Proposition 5], the author proved that the above result also holds for non-principal characters. The present result is consistent with that earlier work.
Earlier, Burgess considered the multidimensional setup for this problem in [5], [6], where
$f(x_{1},x_{2},\ldots,x_{n})=\prod_{i=1}^{n}L_{i}(x_{1},x_{2},\ldots,x_{n})$
is a product of $n$ linear forms over $\mathbb{Z}$, which are linearly independent over $\mathbb{F}_{p}$, with $p$ a prime. He proved that a nontrivial upper bound for the corresponding character sum can be obtained when the interval lengths are all of the same size $H=q^{\rho_{n}+\varepsilon}$. Bourgain and Chang incorporated ideas from Minkowski’s Subspace Theorem and geometry of numbers to get a nontrivial upper bound of the above character sum in a shorter range of $H$ which is $H=q^{\frac{1}{4}+\varepsilon}$, thus obtaining a $n$-dimensional result as strong as Burgess’s original threshold in each dimension (see [2]0. More recently, L.Pierce and J.Xu [19] have established a Burgess type estimate for character sum evaluated on homogeneous forms of arbitrary degrees. In particular, they obtained a nontrivial upper bound for the character sum on the binary admissible forms if lengths of the intervals are of size $p^{\frac{1}{3}+\varepsilon}.$ However, our Theorem 3 provides a nontrivial upper bound of the character sum on the cubic binary if intervals are of size at least $q^{\frac{3}{8}+\varepsilon}$. Although the sizes of the intervals are slightly larger here compared to Pierce’s multidimensional result [19], we derive it by an alternative technique of "Multiplicative Energy of Set Tuples" in the context of character sum and by the use of Theorem [1]. In contrast to Pierce’s and Xu’s Stratification method [19] we shall use simple sum-product techniques to derive a Burgess type estimate for cubic binary forms at a cost of minor increment in the interval lengths. We shall introduce some definitions below to get a flavor of the additive combinatorial techniques used in our study.
Acknowledgments. The author expresses his sincere gratitude to Professor Stephan Baier for bringing this problem to his attention and for his valuable guidance. The author thanks the Ramakrishna Mission Vivekananda Educational and Research Institute for an excellent work environment. The research of the author was supported by a CSIR Ph.D. fellowship under file number 09/0934(13170)/2022-EMR-I.
2. Notations
Throughout this article, we will consider multiplying in the ring $R=\mathbb{F}_{p^{d}}$ (multiplication defined by the multiplication of the field) or $\mathbb{F}_{p}\times\mathbb{F}_{p}$ (multiplication is defined coordinate-wise, $(a,b)\times(c,d)=(ac,bd)$. We shall also denote the group of invertible elements of the ring $R$ by $R^{*}$. Let $\chi$ be a multiplicative character of $R$. Our notation is the following:
  1. (1)
    $AB:=\{ab:a\in A\text{ and }b\in B\},$
  2. (2)
    $aB:=\{a\}B,$
  3. (3)
    The multiplicative energy of $A_{1},A_{2},A_{3},\ldots,A_{n}\subset R$ is defined as
    $E(A_{1},A_{2},\ldots,A_{n}):=|\{(a_{1},\ldots,a_{n},a_{1}^{\prime},\ldots,a_{n}^{\prime}):a_{1}\cdots a_{n}=a_{1}^{\prime}\cdots a_{n}^{\prime}\text{ and }a_{j},a_{j}^{\prime}\in A\cap R^{*}\}|.$
    Using multiplicative characters we can observe:
  4. (4)
    $E(A_{1},A_{2},\ldots,A_{n}):=\frac{1}{|R^{*}|}\sum_{\chi}\prod_{i=i}^{n}\left|\sum_{\eta\in A_{i}}\chi(\eta_{i})\right|^{2}$.
    Note that energy is always multiplicative in this article. Throughout the sequel, $c$ will be an arbitrarily small positive number which, depending on the context, may change from a line to the next. For $n\in\mathbb{N}$, we denote by $d(n)$ the number of divisors of $n$ and recall the well-known bound $d(n)\leq\exp\left(\frac{\log{n}}{\log\log{n}}\right)$.
3. Multiplicative energy of intervals in $\mathbb{F}_{p^{d}}$
First we shall obtain the following lemma which is required for calculating the multiplicative energy of a set triple.
Lemma 4.
Let $\omega\in\mathbb{F}_{p^{d}}$ be a generator over $\mathbb{F}_{p}$. Given $0<\sigma<\tfrac{11}{16}$, define
$Q=\left\{\,x_{0}+x_{1}\omega+x_{2}\omega^{2}+\cdots+x_{d-2}\omega^{d-2}\;:\;x_{i}\in[1,p^{\sigma}]\text{ for all }i\in\{0,1,\ldots,d-2\}\,\right\},$
and
$Q_{0}=\left\{\,y_{0}+y_{1}\omega\;:\;y_{i}\in[1,p^{\frac{9}{20}-\frac{4}{20}\sigma}]\,\right\}.$
Then, for all $\eta\in\mathbb{F}_{p^{d}}$,
$\max_{\eta\in\mathbb{F}_{p^{d}}}\bigl|\{\,(z_{1},z_{2})\in Q\times Q_{0}:\eta=z_{1}z_{2}\,\}\bigr|<\exp\!\left(c_{d}\frac{\log p}{\log\log p}\right),$
where $c_{d}>0$ is a constant depending only on $d$.
Proof.
Fix $\eta\in\mathbb{F}_{p^{d}}$ and suppose that $\eta$ can be represented as a product of two elements, one from $Q$ and one from $Q_{0}$, in at least two distinct ways. That is, assume
$(y_{0}+\omega y_{1}),(y_{0}^{\prime}+\omega y_{1}^{\prime})\in Q_{0},\quad(x_{0}+x_{1}\omega+\cdots+x_{d-2}\omega^{d-2}),(x_{0}^{\prime}+x_{1}^{\prime}\omega+\cdots+x_{d-2}^{\prime}\omega^{d-2})\in Q,$
such that
(1) $(x_{0}+x_{1}\omega+\cdots+x_{d-2}\omega^{d-2})(y_{0}+\omega y_{1})=(x_{0}^{\prime}+x_{1}^{\prime}\omega+\cdots+x_{d-2}^{\prime}\omega^{d-2})(y_{0}^{\prime}+\omega y_{1}^{\prime})=\eta.$
Hence, $\omega$ is a zero of the polynomial
$P(X)=(x_{0}+x_{1}X+\cdots+x_{d-2}X^{d-2})(y_{0}+Xy_{1})-(x_{0}^{\prime}+x_{1}^{\prime}X+\cdots+x_{d-2}^{\prime}X^{d-2})(y_{0}^{\prime}+Xy_{1}^{\prime})$
in $\mathbb{F}_{p}[X]$. Since the minimal polynomial of $\omega$ over $\mathbb{F}_{p}$ has degree $d$, it follows that all coefficients of $P(X)$ must vanish in $\mathbb{F}_{p}$. Moreover, since each coefficient is of magnitude strictly less than $p$, all these equalities also hold over $\mathbb{Z}$.
[…]
(2) $(x_{0}+x_{1}\widehat{\omega}+\cdots+x_{d-2}\widehat{\omega}^{d-2})(y_{0}+\widehat{\omega}y_{1})=(x_{0}^{\prime}+x_{1}^{\prime}\widehat{\omega}+\cdots+x_{d-2}^{\prime}\widehat{\omega}^{d-2})(y_{0}^{\prime}+\widehat{\omega}y_{1}^{\prime}).$
[…]
$\bigl|\{(z_{1},z_{2})\in\widehat{Q}\times\widehat{Q}_{0}:\widehat{\eta}=z_{1}z_{2}\}\bigr|<\exp\!\left(c\frac{\log p}{\log\log p}\right),$
where
$\widehat{Q}=\{\,x_{0}+x_{1}\widehat{\omega}+\cdots+x_{d-2}\widehat{\omega}^{d-2}:x_{i}\in[1,p^{\sigma}]\,\},\quad\widehat{Q}_{0}=\{\,y_{0}+y_{1}\widehat{\omega}:y_{i}\in[1,p^{\frac{9}{20}-\frac{4}{20}\sigma}]\,\}.$
This follows from the divisor bound for algebraic integers in $\mathbb{Z}[\widehat{\omega}$]. Indeed, for $\widehat{\zeta}\in\mathbb{Z}[\widehat{\omega}$],
$d_{Q(\widehat{\omega})}(\widehat{\zeta})^{2}=d\bigl(N_{Q(\widehat{\omega})/\mathbb{Q}}(\widehat{\zeta})\bigr)^{2}\leq\exp\!\left(2\,\frac{\log N_{Q(\widehat{\omega})/\mathbb{Q}}(\widehat{\zeta})}{\log\log N_{Q(\widehat{\omega})/\mathbb{Q}}(\widehat{\zeta})}\right)\leq\exp\!\left(2d^{2}(1+o(1))\,\frac{\log d\,\log p}{\log\log p}\right).$
Here, the middle inequality follows from the classical divisor estimate $d(n)\leq\exp\!\left(\frac{\log n}{\log\log n}\right),$ and the final bound arises from the size of the coefficients defining the elements of $\widehat{Q}$ and $\widehat{Q}_{0}$. ∎
The consequence here will be estimation of multiplicative energy of two intervals in $\mathbb{F}_{p^{d}}$.
Corollary 5.
Let $Q,Q_{0}$ be as in Lemma 4. Then
$E(Q,\,Q_{0})\;\leq\;\exp\!\Bigl(c_{d}\frac{\log p}{\log\log p}\Bigr)\,|Q|\,|Q_{0}|\;,$
Now we require a lemma where we consider the multiplicative energy of a set tuple.
Lemma 6.
Let $Q,Q_{0}$ be as in the previous lemmas and let $I_{s}=[1,p^{1/k_{s}}]$ for $s=1,2,\ldots,r$ with $k_{s}\in\mathbb{Z}_{+}$ and $\sum\frac{1}{k_{s}}<1$. Then
$E(I_{1},I_{2},\ldots,I_{r},Q,Q_{0})\ \leq\exp(c\frac{\log p}{\log\log p})|Q||Q_{0}|\prod_{s}|I_{s}|^{\frac{1}{10}(19-4\sigma)}.$
Proof.
The proof is similar to [8, Lemma 9], but we still give a proof here for the sake of completeness. First we start with the multiplicative energy of intervals:
$E(I_{1},I_{2},\ldots,I_{r},Q,Q_{0})=\frac{1}{p^{d}}\sum_{\chi}\prod\left|\sum_{t\in I_{s}}\chi(t)\right|^{2}\left|\sum_{\eta\in Q}\chi(\eta)\right|^{2}\left|\sum_{\eta\in Q_{0}}\chi(\eta)\right|^{2}.$
Here we shall use Hölder’s inequality to get the following:
(3) $\prod_{s=1}^{r}\left\{\frac{1}{p^{d}}\sum_{\chi}|\sum_{t\in I_{s}}\chi(t)|^{2k_{s}}|\sum_{\eta\in Q}\chi|^{2}|\sum_{\eta\in Q_{0}}\chi|^{2}\right\}^{\frac{1}{k_{s}}}\left\{\frac{1}{p^{d}}\sum_{\chi}|\sum_{\eta\in Q}\chi|^{2}|\sum_{\eta\in Q_{0}}\chi|^{2}\right\}^{1-\sum\frac{1}{k_{s}}}.$
Notation: $\sum_{t\in I_{s}}\chi:=\sum_{t\in I_{s}}\chi(t),\sum_{\eta\in Q}\chi:=\sum_{\eta\in Q}\chi(\eta).$ For simplicity we denote
$A_{s}:=\left\{\frac{1}{p^{d}}\sum_{\chi}|\sum_{t\in I_{s}}\chi(t)|^{2k_{s}}|\sum_{\eta\in Q}\chi|^{2}|\sum_{\eta\in Q_{0}}\chi|^{2}\right\}$
and
$B:=\left\{\frac{1}{p^{d}}\sum_{\chi}|\sum_{\eta\in Q}\chi|^{2}|\sum_{\eta\in Q_{0}}\chi|^{2}\right\}.$
Now by Corrolary 5 we get,
(4) $B=E(Q,Q_{0})\leq\exp\left(c\frac{\log p}{\log\log p}\right)|Q||Q_{0}|.$
It is well known from the definition of multiplicative energy by character sum that
[…]
Let $Q=K+K\omega+\ldots+K\omega^{d-2}$ with $K=[1,p^{\rho}]$. Then
$E(\mathbb{F}_{p},Q)$
$=$ $|\{(t_{1},t_{2},\zeta_{1},\zeta_{2})\in\mathbb{F}_{p}^{2}\times Q^{2}:t_{1}\zeta_{1}=t_{2}\zeta_{2}\neq 0\}|$
$=$ $|\{(t_{1},t_{2},x_{0},\ldots,x_{d-2},x_{0}^{\prime},\ldots,x_{d-2}^{\prime})\in\mathbb{F}_{p}^{2}\times K^{2(d-1)}:t_{1}a_{i}=t_{2}a_{i}\forall i\neq n_{j}\text{ and}$
$t_{1}(x_{0}\omega^{n_{0}}+\ldots+x_{d-2}\omega^{n_{d-2}})=t_{2}(x^{\prime}_{0}\omega^{n_{0}}+\dots+x^{\prime}_{d-2}\omega^{n_{d-2}})$
$(\neq 0)\}|$
$\leq$ $|\{(t_{1},t_{2},x_{0},\ldots,x_{d-2},x_{0}^{\prime},\ldots,x_{d-2}^{\prime})\in\mathbb{F}_{p}^{2}\times K^{2(d-1)}$
$:t_{1}x_{i}=t_{2}x_{i}^{\prime}\neq 0\text{ (for at least one $i$)}\}|$
Equating the coefficient of the above equation we get
[…]
With at least one of the $t_{1}x_{i}=t_{2}x_{i}^{\prime}\neq 0.$ Let us assume that $t_{1}x_{0}=t_{2}x_{0}^{\prime}\neq 0$. Then the above equations (6) becomes
[…]
For i=1, the number of solutions $(x_{0},x_{0}^{\prime},x_{1},x_{1}^{\prime})$ in (7) is bounded by $E([1,p^{\sigma}],[1,p^{\sigma}])$, which is bounded by $p^{2\sigma}\log p$, by [11]. Also the choice of $t\in\mathbb{F}_{p}$ is bounded by $p$ and each of the $x_{2},x_{3},\ldots,x_{d-2}$ is bounded by $p^{\sigma}$. Therefore,
(8) $E(\mathbb{F}_{p},Q)\leq dp^{1+\sigma(d-1)}\log p.$
and using the above estimation in the equation 5 we obtain
(9) $A_{s}\leq|Q_{1}|^{2}\exp(c_{k_{s}}\frac{\log p}{\log\log p}).p^{1+\sigma(d-1)}.$
Let us remind that $|Q|=p^{(d-1)\sigma}$ and $|Q_{0}|=p^{\frac{1}{10}(9-4\sigma)}.$ Combining the equations 3, 4 and 9 we get the following:
$E(I_{1},I_{2},\ldots,I_{r},Q,Q_{0})$
$\leq$ $\exp\left(c\frac{\log p}{\log\log p}\right)|Q_{0}|^{2\sum_{s}1/k_{s}}p^{(1+(d-1)\sigma)\sum_{s}1/k_{s}}(|Q||Q_{0}|)^{1-\sum_{s}1/k_{s}}$
$=$ $\exp\left(c\frac{\log p}{\log\log p}\right)|Q_{0}||Q||Q_{0}|^{\sum_{s}1/k_{s}}|Q|^{-\sum_{s}1/k_{s}}p^{(1+(d-1)\sigma)\sum_{s}1/k_{s}}$
$=$ $\exp\left(c\frac{\log p}{\log\log p}\right)|Q_{0}||Q|p^{\frac{1}{10}(19-4\sigma)\sum_{s}1/k_{s}}$
$=$ $\exp\left(c\frac{\log p}{\log\log p}\right)|Q_{0}||Q|\prod_{s}|I_{s}|^{\frac{1}{10}(19-4\sigma)}$
4. Proof of main theorems
We shall directly prove the general theorem which is Theorem 2.
Proof of Theorem 2:
Proof.
First, we mention that we wish to take $\rho>\frac{3}{8}$. Let
$Q_{0}=\{y_{0}+y_{1}\omega:y_{i}\in[1,p^{\frac{9}{20}-\frac{4}{20}\rho}]\}$
Let us choose $k_{1},\ldots,k_{r}\in\mathbb{N}$ which satisfy
$\frac{6}{5}\rho-\frac{9}{20}-2\varepsilon<\frac{1}{k_{1}}+\ldots+\frac{1}{k_{s}}<\frac{6}{5}\rho-\frac{9}{20}-\varepsilon$
where $\epsilon>0$ can be taken sufficiently small. Let
$I=[1,p^{\varepsilon/2}]\text{ and }I_{s}=[1,p^{1/k_{s}}]$
for $s=1,\ldots,r.$ We then translate $Q$ by
$I.\prod_{s=1}^{r}I_{s}.Q_{0}$
and proceed with the Burgess argument as above. Now it is enough to estimate
$\frac{1}{(|I|.\prod_{s=1}^{r}|I_{s}|.|Q_{0}|)}\sum\limits_{\begin{subarray}{c}q\in Q_{0}\\ (s_{1},\ldots,s_{n})\in\prod_{s}{I_{s}}\\ (x_{0},x_{1},\ldots,x_{d-2})\in\prod_{i=0}^{d-2}J_{i}\end{subarray}}\left|\sum_{t\in I}\chi\left(t+\frac{x_{0}\omega^{n_{0}}+x_{1}\omega^{n_{1}}+\ldots+x_{d-2}\omega^{n_{d-2}}}{s_{1}.s_{2}.\ldots.s_{n}.q}\right)\right|$
Let $\Phi(\mu):=\left|\left\{(x_{0},x_{1},\ldots,x_{d-2},s_{1},\ldots,s_{n},q)\in\prod_{i=0}^{d-2}J_{i}\times\prod_{s=1}^{r}I_{s}\times Q_{0}:\mu=\frac{x_{0}\omega^{n_{0}}+x_{1}\omega^{n_{1}}+\ldots+x_{d-2}\omega^{n_{d-2}}}{s_{1}.\dots.s_{n}.q}\right\}\right|.$ Then the sum can be rephrased as
$\frac{1}{(|I|.\prod_{s=1}^{r}|I_{s}|.|Q_{0}|)}\sum_{\mu\in\mathbb{F}_{p^{d}}}\Phi(\mu)\left|\sum_{t\in I}\chi(t+\mu)\right|.$
Now by using Hölder ineqality twice we get the above sum is bounded by
$\frac{1}{(|I|.\prod_{s=1}^{r}|I_{s}|.|Q_{0}|)}\left(\sum_{\mu\in\mathbb{F}_{p^{d}}}\phi(\mu)^{\frac{2k}{2k-1}}\right)^{1-\frac{1}{2k}}\left(\sum_{\mu\in\mathbb{F}_{p^{d}}}\left|\sum_{t\in I}\chi(\mu+t)\right|^{2k}\right)^{\frac{1}{2k}}$
$\leq\frac{1}{(|I|.\prod_{s=1}^{r}|I_{s}|.|Q_{0}|)}\underbrace{\left[\sum_{\mu\in\mathbb{F}_{p^{d}}}\Phi(\mu)\right]^{1-\frac{1}{k}}}_{\alpha}\underbrace{\left[\sum_{\mu\in\mathbb{F}_{p^{d}}}\Phi(\mu)^{2}\right]^{\frac{1}{2k}}}_{\beta}\underbrace{\left(\sum_{\mu\in\mathbb{F}_{p^{d}}}\left|\sum_{t\in I}\chi(\mu+t)\right|^{2k}\right)^{\frac{1}{2k}}}_{\gamma}$
By trivial estimate we get
(10) $\alpha\leq(|Q||Q_{0}|p^{\sum 1/k_{s}})^{1-\frac{1}{k}}.$
By the definition of multiplicative energy and thereafter by Lemma 6 we get
(11) $\beta=$ $E(Q,Q_{0},I_{1},I_{2},\ldots,I_{r})^{\frac{1}{2k}}$
(12) $\leq$ $|Q||Q_{0}|p^{\frac{1}{10}(19-4\rho)\sum_{s}1/k_{s}}$
By Weil bound [10] we get
(13) $\gamma\leq p^{\frac{\varepsilon}{4}+\frac{d}{2k}}+p^{\frac{\varepsilon}{2}+\frac{d}{4k}}$
CLAIM:
$\beta=|Q||Q_{0}|p^{\frac{1}{10}(19-4\rho)\sum\frac{1}{k_{s}}}<|Q|^{2}|Q_{0}|^{2}p^{2\sum\frac{1}{k_{s}}-\frac{d}{2}-\tau}$
If we assume the above claim for the moment combining the estimates of $\alpha$(10), $\beta$(11) and $\gamma$(13) we get:
$\sum_{q\in Q}\chi(q)$ $\leq\frac{1}{(|I|.\prod_{s=1}^{r}|I_{s}|.|Q_{0}|)}\left(|Q||Q_{0}|p^{\sum_{s}\frac{1}{k_{s}}}\right)^{1-\frac{1}{k}}$
$\left(|Q|^{2}|Q_{0}|^{2}p^{2\sum\frac{1}{k_{s}}-\frac{d}{2}-\tau}\right)^{\frac{1}{2k}}\left(p^{\frac{\varepsilon}{4}+\frac{d}{2k}}+p^{\frac{\varepsilon}{2}+\frac{d}{4k}}\right)$
$\leq\frac{p^{-\frac{\varepsilon}{2}-\sum_{s}\frac{1}{k_{s}}}}{|Q_{0}|}\left(|Q||Q_{0}|p^{\sum_{s}\frac{1}{k_{s}}}\right)^{1-\frac{1}{k}}\left(|Q|^{2}|Q_{0}|^{2}p^{2\sum\frac{1}{k_{s}}-\frac{d}{2}-\tau}\right)^{\frac{1}{2k}}\left(p^{\frac{\varepsilon}{4}+\frac{d}{2k}}+p^{\frac{\varepsilon}{2}+\frac{d}{4k}}\right)$
$\leq|Q|\left(p^{\frac{-\varepsilon}{4}+\frac{1}{2k}(\frac{d}{2}-\tau)}+p^{-\frac{\tau}{2k}}\right)$
Now if we take $k$ to sufficiently large precisely $k>\frac{d}{\varepsilon}$ then Theorem 2 is proved.
Proof of claim. It is enough to show
$(d-1)\rho+(\frac{9}{10}-\frac{4}{10}\rho)+\frac{1}{10}(19-4\rho)\sum_{s}\frac{1}{k_{s}}<2(d-1)\rho+2(\frac{9}{10}-\frac{4}{10}\rho)+2\sum_{s}\frac{1}{k_{s}}-\frac{d}{2}$
This is equivalent to
$(d-1)\rho+(\frac{9}{10}-\frac{4}{10}\rho)+\frac{1}{10}(4\rho+1)\sum_{s}\frac{1}{k_{s}}-\frac{d}{2}>0$
From the choices of $k_{s}$ in the initial steps we get
$(d-1)\rho+(\frac{9}{10}-\frac{4}{10}\rho)+\frac{1}{10}(4\rho+1)(\frac{6}{5}\rho-\frac{9}{20})-\frac{d}{2}>0$
By simplifying this we get
$480\rho^{2}+(1000d-1460)\rho+(855-500d)>0$
$\iff\rho>\frac{1460-1000d+\sqrt{1000000d^{2}-1960000d+490000}}{960}$
Note: $\rho_{3}=3/8=.375,\rho_{4}=.417,\rho_{5}=.438,\rho_{6}=.447\text{ and so on.}$ Also, we can observe that $\rho_{d}<1/2$. So, it is indeed a short-character sum over intervals.
Proof of Theorem 1 comes as a corollary of Theorem 2 if we take $d=3.$
5. Application to character sums of binary cubic form
First we would like to illustrate an application of our theorem:
Theorem.
Given $\varepsilon>0$, there is $\delta>0$ such that the following holds. Let $p$ be a large prime and $f(x,y)=x^{3}+ax^{2}y+bxy^{2}+cy^{3}$ is non-degenerate over $\mathbb{F}_{p}$. Let $I,J\subseteq[1,p-1]$ be interval of size
$p^{\frac{1}{2}}>|I|,|J|>p^{\frac{3}{8}+\varepsilon}$
Then
$\left|\sum_{x\in I,y\in J}\chi(f(x,y))\right|<p^{-\delta}|I||J|$
for any non principle multiplicative character $\chi(\bmod{p)}$. This estimate is uniform in $f$.
A similar result was shown by Burgess assuming $|I|,|J|>p^{\frac{1}{3}}$ and consequently by Chang assuming $|I|,|J|>p^{\frac{1}{4}}$ for the binary quadratic forms. But here we consider binary cubic forms.
Proof.
We now distinguish three non-degenarate cases for $f(x,y)$. Here by non-degenerate cubic form we mean that the de-homogenized polynomial of $f(x,y)$ cannot have a root of multiplicity greater than $1$.
Case 1. $f$ is irreducible $(\bmod{p})$. Then $\chi(f(x,y))$ is a character $\bmod{p}$ of $x+\omega y$ with $\omega$ is a root of the polynomial $g(x):=x^{3}+ax^{2}+bx^{2}+cx+d$ in the cubic extension $Q(\omega)$ and the result then follows from Theorem 1.
Case 2. $f(x,y)$ is reducible in $\mathbb{F}_{p}[x,y]$ where
$f(x,y)=(x-\lambda_{1}y)(x^{2}-\lambda_{2}xy+\lambda_{3}y^{2})=(x-\lambda_{1}y)(x-\omega_{2}y)(x-\omega_{3}y).$
where $\lambda_{1}\in\mathbb{F}_{p}$ and $\omega_{2},\omega_{3}\in\mathbb{F}_{p^{2}}$ We shall proceed similar to M.C. Chang [Theorem 11,[8]] but replacing $\mathbb{F}_{p}\times\mathbb{F}_{p}$ by $\mathbb{F}_{p}\times\mathbb{F}_{p^{2}}\times\mathbb{F}_{p^{2}}.$
Denote:
$I_{0}=[1,\frac{1}{4}p^{3/8}]$ and $K=[1,p^{\kappa}]$. We translate $(x,y)$ by $(stx_{0},sty_{0})$ with $x_{0},y_{0}\in I_{0}$ and $s,t\in K$ and estimate
(14) $\frac{1}{|K|^{2}|I_{0}|^{2}}\sum\limits_{\begin{subarray}{c}x\in I,y\in J\\ x_{0},y_{0}\in I_{0}\\ s\in K\end{subarray}}\left|\sum_{t\in K}\chi\left(\left(t+\frac{x-\lambda_{1}y}{s(x_{0}-\lambda_{1}y_{0})}\right)\left(t+\frac{x-\omega_{2}y}{s(x_{0}-\omega_{2}y_{0})}\right)\left(t+\frac{x-\omega_{3}y}{s(x_{0}-\omega_{3}y_{0})}\right)\right)\right|$
So here if we denote
(15) $w(z_{1},z_{2}):=\\ \left|\left\{(x,y,x_{0},y_{0},s)\in I\times J\times I_{0}\times I_{0}\times K:z_{1}=\frac{x-\lambda_{1}y}{s(x_{0}-\lambda_{1}y_{0})},z_{2}=\frac{x-\omega_{2}y}{s(x_{0}-\omega_{2}y_{0})}\right\}\right|.\\$
By this notation the above estimate will become
(16) $\frac{1}{|K|^{2}|I_{0}|^{2}}\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p^{2}}\end{subarray}}w(z_{1},z_{2})\left|\sum_{t\in K}\chi((t+z_{1})(t+z_{2})(t+z_{2}^{\prime}))\right|\\$
where $z_{2},z_{2}^{\prime}$ are the roots of a monic polynomial over $\mathbb{F}_{p}$. We shall use Holder’s inequality subsequently Weil’s Theorem to estimate the above. Using Holder’s inequality twice it is obtained
(17) $\frac{1}{|K|^{2}|I_{0}|^{2}}\left(\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p^{2}}\end{subarray}}w(z_{1},z_{2})\right)^{1-\frac{1}{k}}\left(\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p^{2}}\end{subarray}}w(z_{1},z_{2})^{2}\right)^{\frac{1}{2k}}\left(\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p^{2}}\end{subarray}}\left|\sum_{t\in K}\chi((t+z_{1})(t+z_{2})(t+z_{2}^{\prime})\right|^{2k}\right)^{\frac{1}{2k}}$
First we bound the first term of the product trivially by
(18) $\left(\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p^{2}}\end{subarray}}w(z_{1},z_{2})\right)^{1-1/k}\leq(|I||J||I_{0}|^{2}|K|)^{1-1/k}$
The third term is estimated using Weil’s bound [9] and consequently by Lemma 11 as the following
$\left(\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p^{2}}\end{subarray}}\left|\sum_{t\in K}\chi((t+z_{1})(t+z_{2})(t+z_{2}^{\prime}))\right|^{2k}\right)^{\frac{1}{2k}}\ll\left(\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ a\in\mathbb{F}_{p}\\ b\in\mathbb{F}_{p}\end{subarray}}\left|\sum_{t\in K}\chi((t+z_{1})(t^{2}+a\,t+b))\right|^{2k}\right)^{\frac{1}{2k}}$
(19) $\ll r|K|^{\frac{1}{2}}p^{\frac{3}{2r}}+|K|p^{\frac{3}{4r}}\\$
The required bound for the second term is the following:
(20) $\sum w(z_{1},z_{2})^{2}\leq|I|^{2}|J|^{2}|K|^{2}p^{-\tau}$
for some $\tau>0$. Choosing $\kappa=3/2k$, we obtain an upper bound for the equation (14) of the form $|I||J|p^{-\tau/2k}$ from the above three estimates. For estimating second term we may remind the reader $|I|,|J|<p^{1/2}$ here. Let
$R=\{(x-\lambda_{1}y,x-\omega_{2}y,x-\omega_{3}y):x\in I,y\in J\}$
$T=\{(x_{1}-\lambda_{1}y_{1},x_{1}-\omega_{2}y_{1},x_{1}-\omega_{3}y_{1}):x_{1},y_{1}\in I_{0}\}$
$S=\{(s,s,s):s\in K\}.$
We can check from the above description
$\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p^{2}}\end{subarray}}w(z_{1},z_{2})^{2}\leq E(R,T,S).$
Now it suffices to show
(21) $E(R,T,S)\leq|I|^{2}|J|^{2}|K|^{2}p^{-\tau}.$
First we estimate the following:
Lemma 7.
Let $R$ and $T$ be defined as above. Then
$E(R,T)<\exp\left(c\frac{\log p}{\log\log p}\right).$
We shall prove the above lemma in due course. Let us start with writing $R$ as union of translates of a $T$ as bellow:
$R=\bigcup_{\alpha<|R|/|T|}(T+\zeta_{\alpha})$
Now we have
$E(R,T)\leq\frac{|R|^{2}}{|T|^{2}}\max_{\zeta\in R}E(T+\zeta,T).$
Now, it is enough to show
$\max_{T+\zeta\,\subseteq\,R}E(T+\zeta,T)<\exp\left(c\frac{\log p}{\log\log p}\right)|T|^{2}.$
This is an immediate consequence of the Lemma below.
Lemma 8.
Let $T$ is given in above. Then
$\max_{\rho\in\mathbb{F}_{p}^{*}\times\mathbb{F}_{p^{2}}^{*}\times\mathbb{F}_{p^{2}}^{*}}\left|\left\{(z_{1}+\zeta_{,}z_{2})\in T+\zeta\times T:\rho=z_{1}z_{2}\right\}\right|<\exp\left(c\frac{\log p}{\log\log p}\right)$
where, $T+\zeta\subseteq R$.
Proof.
Let us fix an $\rho=(\rho_{1},\rho_{2},\rho_{3})\in\mathbb{F}_{p}^{*}\times\mathbb{F}_{p^{2}}^{*}\times\mathbb{F}_{p^{2}}^{*}$ and consider a product representation of $\rho=(z_{1}+\zeta)\times z_{2}$ where $z_{1},z_{2}\in T$. Then $z_{1}=(x_{1}-\lambda_{1}y_{1},x_{1}-\omega_{2}y_{1},x_{1}-\omega_{3}y_{1})$, $z_{2}=(x_{2}-\lambda_{1}y_{2},x_{2}-\omega_{2}y_{2},x_{2}-\omega_{3}y_{2})$ and $\zeta=(x_{\zeta}-\lambda_{1}y_{\zeta},x_{\zeta}-\omega_{2}y_{\zeta},x_{\zeta}-\omega_{3}y_{\zeta})$ with $x_{1},x_{2},y_{1},y_{2}\in I_{0},x_{\zeta}\in I$ and $y_{\zeta}\in J$, we want to obtain an estimate of the number of solutions of the following:
(22) $\begin{cases}(x_{1}+x_{\zeta}-\lambda_{1}(y_{1}+y_{\zeta}))(x_{2}-\lambda_{1}y_{2})=\rho_{1}(\text{ in }F_{p}(\omega))\\ \\ (x_{1}+x_{\zeta}-\omega_{2}(y_{1}+y_{\zeta}))(x_{2}-\omega_{2}y_{2})=\rho_{2}(\text{ in }F_{p}(\omega))\\ \\ (x_{1}+x_{\zeta}-\omega_{3}(y_{1}+y_{\zeta}))(x_{2}-\omega_{3}y_{2})=\rho_{3}(\text{ in }F_{p}(\omega))\end{cases}$
where $\omega_{2},\omega_{3}$ are the root of the polynomial $x^{2}+\lambda_{2}x+\lambda_{3}.$ The estimate of number of solutions of (22) is again bounded by the number of solution of the following:
(23) $\begin{cases}(x_{1}-\lambda_{1}y_{1})(x_{2}-\lambda_{1}y_{2})=\rho_{1}(\text{ in }F_{p}(\omega))\\ \\ (x_{1}-\omega_{2}y_{1})(x_{2}-\omega_{2}y_{2})=\rho_{2}(\text{ in }F_{p}(\omega))\\ \\ (x_{1}-\omega_{3}y_{1})(x_{2}-\omega_{3}y_{2})=\rho_{3}(\text{ in }F_{p}(\omega))\end{cases}$
where, $x_{1}\in I,y_{1}\in J$ and $x_{2},y_{2}\in I_{0}$.
Let $\mathcal{S}$ be the solution set of the above equation. Let $(x_{1},x_{2},y_{1},y_{2}),(x_{1}^{\prime},x_{2}^{\prime},y_{1}^{\prime},y_{2}^{\prime})\in\mathcal{S}$. So, it satisfies the equation (22). So, all three of $\lambda_{1},\omega_{1},\omega_{3}$ satisfies the quadratic polynomial equation
$(y_{1}y_{2}-y_{1}^{\prime}y_{2}^{\prime})X^{2}+(x_{1}^{\prime}y_{2}^{\prime}+y_{1}^{\prime}x_{2}-x_{1}y_{2}-y_{1}x_{2})X+(x_{1}x_{2}-x_{1}^{\prime}x_{2}^{\prime})=0\text{ in }\mathbb{F}_{p^{2}}.$
It implies that
[…]
Also in the above congruences (24) the elements are in $\mathbb{F}_{p}$. So, this gives the same congruences in the field $\mathbb{F}_{p}$. Now by the definition of $I,J$ and $I_{0}$; their elements are bounded by $p^{1/2}$ and $\frac{p^{3/8}}{4}$ respectively. Therefore, $|x_{1}x_{2}-x_{1}^{\prime}x_{2}^{\prime}|,|y_{1}y_{2}-y_{1}^{\prime}y_{2}^{\prime}|$ are bounded by $p^{7/8}$. As both of them are elements of $\mathbb{Z}$ with an absolute value less than the modulus the equalities in (24) are, in fact, inequalities in $\mathbb{Z}$ with
$x_{1}x_{2}=x_{1}^{\prime}x_{2}^{\prime}$
and
$y_{1}y_{2}=y_{1}^{\prime}y_{2}^{\prime}.$
Finally we shall again use the divisor estimate $\exp\left(c\frac{\log p}{\log\log p}\right)$ to get the required result. ∎
Now we shall continue with proving the equation (21). Let $\kappa=3/2k$ in the definition of $K$. Thus,
$E(R,T,S)$
$=$ $\frac{1}{p^{2}}\sum_{\chi=\chi_{1}\chi_{2}\chi_{3}}\left|\sum_{z\in S}\chi(z)\right|^{2}\left|\sum_{z_{1}\in R}\chi(z_{1})\right|^{2}\left|\sum_{z_{2}\in T}\chi(z_{2})\right|^{2}$
$\leq$ $\left[\frac{1}{p^{2}}\sum_{\chi}\left|\sum_{z\in S}\chi(z)\right|^{2k}\left|\sum_{R}\cdots\right|^{2}\left|\sum_{T}\cdots\right|^{2}\right]^{\frac{1}{k}}\left[\frac{1}{p^{2}}\sum_{\chi}\left|\sum_{R}\cdots\right|^{2}\left|\sum_{T}\cdots\right|^{2}\right]^{1-\frac{1}{k}}$
$=$ $Q^{\frac{1}{k}}E(R,T)^{1-\frac{1}{k}}$
[…]
(25) $E(R,T,S)\leq Q^{\frac{1}{k}}.\exp\left(c\frac{\log p}{\log\log p}\right).|R|^{2(1-\frac{1}{k})}$
where, $Q=\left[\frac{1}{p^{2}}\sum_{\chi}\left|\sum_{z\in S}\chi(z)\right|^{2k}\left|\sum_{R}\cdots\right|^{2}\left|\sum_{T}\cdots\right|^{2}\right]^{\frac{1}{k}}$
$Q=$ $\frac{1}{p^{2}}\sum_{\chi}\left|\sum_{z\in S}\chi(z)\right|^{2k}\left|\sum_{R}\cdots\right|^{2}\left|\sum_{T}\cdots\right|^{2}$
$\leq$ $\frac{|T|^{2}}{p^{2}}\sum_{\chi}\left|\sum_{z\in S}\chi(z)\right|^{2k}\left|\sum_{z_{1}\in R}\chi(z_{1})\right|^{2}$
$<$ $\exp\left(c_{k}\frac{\log p}{\log\log p}\right).\frac{|T|^{2}}{p^{2}}\sum_{\chi_{1},\chi_{2},\chi_{3}}\left|\chi_{1}(t)\chi_{2}(t)\chi_{3}(t)\right|^{2}\left|\sum_{x\in I,y\in J}\chi_{1}(x-\lambda_{1}y)\chi_{2}(x-\omega_{2}y)\chi_{3}(x-\omega_{3}y)\right|^{2}$
$=$ $\exp\left(c.\frac{\log p}{\log\log p}\right).|T|^{2}E(R,\Delta)$
So,
(26) $Q\leq\exp\left(c.\frac{\log p}{\log\log p}\right).|T|^{2}E(R,\Delta)$
where $\Delta:=\{(t,t,t):t\in\mathbb{F}_{p}\}.$ The multiplicative energy $E(R,\Delta)$ above evaluates the number of solutions $(x,x^{\prime},y,y^{\prime},t,t^{\prime})\in I^{2}\times J^{2}\times(\mathbb{F}^{*}_{p})^{2}$ of the following simultaneous linear congruent equations
$(x-\lambda_{1}y,x-\omega_{2}y,x-\omega_{3})(t,t,t)=(x^{\prime}-\lambda_{1}y^{\prime},x^{\prime}-\omega_{2}y^{\prime},x^{\prime}-\omega_{3}y^{\prime})(t^{\prime},t^{\prime},t^{\prime}).$
It gives the following equations:
[…]
From the above it follows that
[…]
which implies $xy^{\prime}=x^{\prime}y\bmod{p}$ and number of solution set of it is bounded by $E(I,J)\lesssim(\log p).|I||J|$ where $|I|,|J|<p$. Once $x.x^{\prime},y,y^{\prime}$ are fixed, the number of solutions $(t,t^{\prime})$ in(27) is at most (p1)(p-1). Hence, the number of solutions of equation(27) is bounded by $p(\log p)|I||J|$. So,
$E(R,\Delta)\lesssim p(\log p).|I|.|J|.$
Using this in 26 we get
(29) $Q\leq\exp\left(c\frac{\log p}{\log\log p}\right)p(\log p)|T|^{2}.|R|.$
Substituting this from (29) into (27), we get
(30) $E(R,T,S)<\exp(c\frac{\log p}{\log\log p})p^{1/k}|R|^{2-1/k}|T|^{2/k}$
From the definition of $T$ we have $|T|=|I_{0}|^{2}=p^{3/4}$. Now by our choice $\kappa=3/2k$ and $|K|=p^{3/2k}$. Substituting the value of $|T|$ in X we obtain
$E(R,T,S)$ $\leq\exp\left(c\frac{\log p}{\log\log p}\right)p^{5/2k}R^{2-1/k}$
$\leq\exp\left(c\frac{\log p}{\log\log p}\right)p^{3/k}(|I||J|)^{2-1/k}$
$\leq\exp\left(c\frac{\log p}{\log\log p}\right)|K|^{2}(|I||J|)^{2-2\kappa/3}$
This proves the above estimate (21).
Case 3. $f(x,y)$ splits completely in $\mathbb{F}_{p}$ with distinct linear factors. Suppose
$f(x,y)=(x-\lambda_{1}y)(x-\lambda_{2}y)(x-\lambda_{3}y).$
The same argument applies over $\mathbb{F}_{p}\times\mathbb{F}_{p}\times\mathbb{F}_{p}$ in place of $\mathbb{F}_{p}\times\mathbb{F}_{p^{2}}$. Here $I_{0}$ and $K$ are the same as before. Translating $(x,y)$ by $(stx_{0},sty_{0})$ we obtain a character sum estimate
(31) $\frac{1}{|K|^{2}|I_{0}|^{2}}\sum\limits_{\begin{subarray}{c}x\in I,y\in J\\ x_{0},y_{0}\in I_{0}\\ s\in K\end{subarray}}\left|\sum_{t\in K}\chi\left(\left(t+\frac{x-\lambda_{1}y}{s(x_{0}-\lambda_{1}y_{0})}\right)\left(t+\frac{x-\lambda_{1}y}{s(x_{0}-\lambda_{2}y_{0})}\right)\left(t+\frac{x-\lambda_{3}y}{s(x_{0}-\lambda_{1}y_{0})}\right)\right)\right|.$
So, here if we denote
$w(z_{1},z_{2},z_{3}):=$
$\left|\left\{(x,y,x_{0},y_{0},s)\in I\times J\times I_{0}\times I_{0}\times K:z_{1}=\frac{x-\lambda_{1}y}{s(x_{0}-\lambda_{1}y_{0})},z_{2}=\frac{x-\lambda_{1}y}{s(x_{0}-\lambda_{1}y_{0})},z_{3}=\frac{x-\lambda_{1}y}{s(x_{0}-\lambda_{1}y_{0})}\right\}\right|$
the above equation becomes
(32) $\frac{1}{|K|^{2}|I_{0}|^{2}}\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p}\\ z_{3}\in\mathbb{F}_{p}\end{subarray}}w(z_{1},z_{2},z_{3})\left|\sum_{t\in K}\chi((t+z_{1})(t+z_{2})(t+z_{3})\right|.$
rf
After applying Hölder’s inequality and the Weil bound (a variant of Lemma 11), it remains to show that
$E(R,T,S)\leq p^{-\tau}|I|^{2}|J|^{2}|K|^{2},$
where
$R$ $=\{(x-\lambda_{1}y,\,x-\lambda_{2}y,\,x-\lambda_{3}y):x\in I,\,y\in J\},$
$T$ $=\{(x_{0}-\lambda_{1}y_{0},\,x_{0}-\lambda_{2}y_{0},\,x_{0}-\lambda_{3}y_{0}):x_{0},y_{0}\in I_{0}\},$
$S$ $=\{(s,s,s):s\in K\}.$
The estimation of $E(R,T,S)$ proceeds exactly as in Case 2, yielding the same bound as required. ∎
Appendix
We recall the classical Weil bound for multiplicative character sums, which plays a key role in our estimates.
Theorem 9 (Weil).
Let $\mathbb{F}$ be a finite field with $q$ elements and let $\chi$ a nontrivial multiplicative character of $\mathbb{F}^{*}$ of order $d>1$. Suppose $f\in\mathbb{F}[X]$ has $m$-distinct roots and $f$ is not a $d$-th power. Then for $n\geq 1$ we have
$\left|\sum_{x\in\mathbb{F}_{n}}\chi(N(f(x)))\right|\leq(m-1)q^{n/2}.$
where, $\mathbb{F}_{n}$ is the degree $n$ extension of $\mathbb{F}$ and $N()$ is the norm.
Proof.
See [16, Chapter 11]. ∎
A more general theorem in this regard is the following
Theorem 10 (Weil bound for multiplicative characters on $\mathbb{F}_{p^{d}}$).
Let $q=p^{n}$ be a power of a prime and let $\chi:\mathbb{F}_{q}^{\times}\to\mathbb{C}^{\times}$ be a nontrivial multiplicative character of order $d$. Extend $\chi$ to $\mathbb{F}_{q}$ by defining $\chi(0)=0$. Let $f(x)\in\mathbb{F}_{q}[x]$ be a polynomial which has $m\geq 1$ distinct root such that and is not an $d$-th power in $\overline{\mathbb{F}}_{q}[x]$. Then
$\left|\sum_{x\in\mathbb{F}_{q}}\chi\big(f(x)\big)\right|\leq(m-1)q^{1/2}.$
Proof.
See the proof of [20, Theorem 2C’, p.43].
The following lemma provides the key averaged estimate used in the main argument.
Lemma 11.
Let $\chi$ be a nontrivial multiplicative character of $\mathbb{F}_{p}^{*}$ of order $d>1$. Then for any integer $r\geq 1$,
$\left(\sum_{a,b,c\in\mathbb{F}_{p}}\Big|\sum_{t\in K}\chi\big((t+a)(t^{2}+bt+c)\big)\Big|^{2r}\right)^{\!\!\frac{1}{2r}}\ll r|K|^{1/2}p^{\frac{3}{2r}}+|K|p^{\frac{3}{4r}}.$
Proof.
Expanding the $2r$-th power, we obtain
$\sum_{a,b,c}\!\Big|\sum_{t\in K}\chi((t+a)(t^{2}+bt+c))\Big|^{2r}=\!\sum_{t_{1},\dots,t_{2r}\in K}\!\sum_{a,b,c}\prod_{i=1}^{2r}\chi_{i}\!\big((t_{i}+a)(t_{i}^{2}+bt_{i}+c)\big),$
where $\chi_{i}=\chi$ for $1\leq i\leq r$ and $\chi_{i}=\overline{\chi}$ for $r<i\leq 2r$. The inner sum factors into two character sums—one over $\mathbb{F}_{p}$ and one over its quadratic extension $\mathbb{F}_{p^{2}}$ as follows
$\sum_{t_{1},t_{2},\ldots,t_{2r}\in K}\left|\sum_{a\in\mathbb{F}_{p}}\chi\left(\prod_{i=1}^{r}(t_{i}+a)\prod_{i=r+1}^{2r}(t_{i}+a)^{p-2}\right)\right|\times\left|\sum_{\omega\in\mathbb{F}_{p^{2}}}\chi\left(N\left(\prod_{i=1}^{r}(t_{i}+\omega)\prod_{i=r+1}^{2r}(t_{i}+\omega)^{p-2}\right)\right)\right|.$
Step 1. Counting exceptional tuples. We call a tuple $(t_{1},\dots,t_{2r})\in K^{2r}$ bad if the polynomial inside the character becomes a $d$-th power in $\mathbb{F}_{p}[X]$; in this case Weil’s bound does not apply. Now in any bad tuple every element occurs at least twice. Consequently, the number of such bad tuples is bounded by
$\{\text{bad tuples}\}\leq|K|^{r}.$
For these tuples, we use the trivial estimate for the character sums, giving a contribution $\leq|K|^{r}p^{3}$.
Step 2. Applying Weil’s bound. For all other (good) tuples, the conditions of Theorem 9 are satisfied, and each of the two character sums are bounded by $O(r\,p^{1/2})$ and O(rp)O(r\,p) respectively. Hence their combined contribution is $\ll r^{2}|K|^{2r}p^{3/2}$.
Step 3. Combining both cases. Summing these contributions we obtain
$\sum_{a,b,c}\Big|\sum_{t\in K}\chi((t+a)(t^{2}+bt+c))\Big|^{2r}\ll|K|^{r}p^{3}+r^{2}|K|^{2r}p^{3/2}.$
Taking the $2r$-th root yields the desired bound. All implied constants are absolute. ∎
Remark 3. Other Weil type estimates of this paper can be proved similarly.